Cusic Industries had the following operating results for last year; sales = $25,700; cost of goods...

  1. Cusic Industries had the following operating results for last year; sales = $25,700; cost of goods sold = $18,400; depreciation expense = $3,450; interest expense = $790. At the beginning of last year, net fixed assets were $19,280, current assets were $5,100, and current liabilities were $3,400. At the end of last year, net fixed assets were $21,650, current assets were $5,830 and current liabilities were $3,580. The tax rate for last year was 40%. Assume there is no short term investments and notes payable.

a. Construct the Income Statement for Cusic Industries. What was net income for last year?

b. What was NOPAT for last year? What caused the difference between net income and NOPAT?

c. What was Return on Invested Capital?

d.   What was the free cash flow for last year? Is this possible? Explain.

Please show all the work in Excel
Answer #1
a] Sales $            25,700
-Cost of goods sold $            18,400
-Depreciation $              3,450
=EBIT $              3,850
-Interest $ 790
=EBT $              3,060
-Tax at 40% $              1,224
=NI $              1,836
b] NOPAT = EBIT*(1-t) = 3850*(1-40%) = $              2,310
The difference is the after tax interest
expense. It is 790*(1-40%) = $ 474
c] Invested capital = Current assets+Net fixed assets-Current liabilities = 5100+19280-3400 = $            20,980
ROIC = NOPAT/Invested capital = 2310/20980 = 11.01%
d] FCF = OCF-Capital expenditure-Changes in NWC
OCF = EBIT*(1-t)+Depreciation = 3850*(1-40%)+3450 = $              5,760
Capital expenditure = Ending NFA-Beginning NFA+Depreciation = 21650-19280+3450 = $              5,820
Changes in NWC = Ending NWC-Beginning NWC = (5830-3580)-(5100-3400) = $ 550
FCF = 5760-5820-550 = $                -610
Negative FCF is possible as the expenditure in FA
and the increase in NWC put together can be more
than OCF.